hopesanddreams3453:

10(1 - 2y) = -5(2y - 1)

Vocaloid:

distribute on both sides, let me know what you get

hopesanddreams3453:

10*1 - 10*2y= -5* 2y- -5*1

Vocaloid:

good, keep simplifying

hopesanddreams3453:

10-20y= -10y- -6 10-6= -20y -10y 4/-10= 10/10 y-0.4

hopesanddreams3453:

are you still there

Vocaloid:

10*1 - 10*2y= -5* 2y- -5*1 10 - 20y = -10y + 5

Vocaloid:

not sure where you got 6 from

Vocaloid:

also, please write - - as + every time you see it

Vocaloid:

so now we have: 10 - 20y = -10y + 5 add 20y to both sides

hopesanddreams3453:

10= -10y= 30y +5

Vocaloid:

check your signs

Vocaloid:

10 - 20y + 20y = ?

hopesanddreams3453:

do i add 20 agian? or 10

Vocaloid:

you start with 10 you take away 20y you put back 20y what's left?

hopesanddreams3453:

-10 + 20y

Vocaloid:

you start with 3 apples you take away 2 apples you put back 2 apples you are back where you started, 3 apples

Vocaloid:

now apply the same logic to the problem

Vocaloid:

you start with 10 you take away 20y you put back 20y what's left?

hopesanddreams3453:

like this: 10-20y+20

hopesanddreams3453:

20y

Vocaloid:

yes, and that equals...?

Vocaloid:

10-20y+20y = ?

hopesanddreams3453:

10-40y

Vocaloid:

hint: A-B+B = A follow the pattern. 10-20y+20y = ?

Vocaloid:

think about the apples problem

hopesanddreams3453:

10-20y=20y

Vocaloid:

you start with 3, take away 2, put back 2, and you have 3 again you start with 10, take away 20y, put back 20y, you have ___ again fill in the blank

hopesanddreams3453:

10-20y=20y

Vocaloid:

nope

Vocaloid:



Vocaloid:

I have drawn an arrow to what goes in the blank

hopesanddreams3453:

10-20y=10

Vocaloid:

almost

Vocaloid:

so, 10 - 20y + 20y = 10 make sense?

hopesanddreams3453:

now i get it

Vocaloid:

so now let's add 20y to the other side

Vocaloid:

-10y + 5+ 20y = ?

hopesanddreams3453:

10=5

hopesanddreams3453:

10+5

Vocaloid:

don't forget the y

Vocaloid:

10y + 5

hopesanddreams3453:

ok

hopesanddreams3453:

is that my answer

Vocaloid:

now we have: 10 = 10y + 5 y = ?

hopesanddreams3453:

sub 5 on both side correct

Vocaloid:

good

hopesanddreams3453:

5=10y

Vocaloid:

hold on I think I made a mistake somewhere

hopesanddreams3453:

ok

Vocaloid:

nvm we're good

hopesanddreams3453:

ok

Vocaloid:

so if 5=10y then y = ?

hopesanddreams3453:

y=0.5

Vocaloid:

good that's your answer